A 300 g bird flying along at 5.5 m/s sees a 10g insect heading straight toward it with a speed of 25 m/s (as measured by an observer on the ground, not by the bird). The bird opens its mouth wide and enjoys a nice lunch.
What is the bird's speed immediately after swallowing?
_____m/s
Answer #1

By conservation of momentum

m1*v1 + m2*v2 = m3*v3

=> 0.3*5.5 + 0.01*(-25) = 0.31*v3

v3 = 4.52 m/s